# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v1

@@ 1. Introduction
For the type $A$ case, the extra structure that we need to use is known as quantum Schur-Weyl duality between the vector representation and the natural representation of the braid group between the tensor powers of the vector representations.
In the classical case, Schur-Weyl duality is a key property for the general linear groups starting with the work of Schur and Weyl. Schur-Weyl duality also takes an important role in Doplicher-Roberts abstract duality theory for compact groups and applications to construction of the field algebra. This classical duality emergences as the starting point for the construction of global gauge symmetry groups in tensor categories arising in [22].
@@
In our work, the need of quantum duality between the braid group representation and the fundamental representation $V$ emerged on the occasion of the proof of our earlier Lemma 7.10, which was not complete in the version 7 of [11], and we worked on completion with the same methods. It was for us a sudden realization of an even greater recognition than we had realized up to that point of the contribution of the authors on extensions of Schur-Weyl duality which has historical interest on his own in representation theory, and more work by Wenzl and his collaborators, and several other authors in the classical and quantum cases for the various Lie types. We shall come back to this in more detail.
@@ 2. The main result
Let ${\mathfrak{g}}$ be a complex simple Lie algebra and let ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ be the module category of the corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ at level $k\in{\mathbb{N}}$, endowed with the ribbon braided tensor structure introduced by Huang and Lepowsky and let $A(V_{{\mathfrak{g}}_{k}})$ be the Zhu algebra. Then
@@
    The representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ of the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ admits a natural structure of ribbon braided tensor category obtained by transporting Huang-Lepowsky structure to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$. In this way, Zhu’s linear equivalence
@@
    Let $A(V_{{\mathfrak{g}}_{k}})$ be endowed with the unitary compatible coboundary weak quasi-Hopf structure induced by ${\mathcal{C}}({\mathfrak{g}},q)$ via the construction of a weak tensor structure on Wenzl functor and of the Drinfeld twist following Theorem 31.1 in [11] (a), and let ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
    Then ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way. Moreover the transport in (a) preserves the tensor products, the ribbon for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples $(V_{\lambda},V,V)$, $(V,V_{\lambda},V)$,$(V,V,V_{\lambda})$ with $V_{\lambda}$ an arbitrary irreducible of the fusion category.
@@
The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary coboundary weak Hopf algebra, are given in Sects. 29, 30 in [11], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [87] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix. The former more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
@@
The mathematical contribution by Kohno, Tsuchiya and Kanie [84], and the subsequent works by A. Wassermann [86] and Toledano-Laredo [83] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [87] that led to our approach. The reader may find reference to more of the several original papers in mathematics and physics in Sections 31–36 in [11].
@@
The centralizer algebra of the fundamental representation for types $B$ and $D$ is studied in [81], [89].
@@ 13. Outlook
 [^1] The claims, arguments, and reasoning of the recent papers [6], [7], [8] are essentially dependent on the results and in-depth discussions between S.C. and C.P. that took place from 2018 to 2022 on the contents of an entire project including [11] and the topics of the above papers, during a period in which S.C. was a co-author, before breaking the collaboration in early 2023, with a minor contribution in a draft of a single section. In addition, C.P. department digital devices were subject to a serious hacking attack for several months in 2021 during the previous collaboration, which was recognized by authorities, that severely damaged her intellectual property and health.
Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of affine vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem. She is also grateful to Marco V. Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [11] and to Hans Wenzl for kind email correspondence and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper. Finally she thanks Pavel Etingof for informing her of [28].

### v2

@@ 1. Introduction
As above mentioned, in the type $A$ case, the extra structure that we need to use is known as quantum Schur-Weyl duality between the vector representation and the natural representation of the braid group between the tensor powers of the vector representations.
In the classical case, Schur-Weyl duality is a key property for the general linear groups starting with the work of Schur and Weyl. We note that Schur-Weyl duality also takes an important role in Doplicher-Roberts abstract duality theory for compact groups and applications to construction of the field algebra. This classical duality emergences as the starting point for the construction of global gauge symmetry groups in tensor categories arising in [22].
@@
In our work, the need of quantum duality between the braid group representation and the fundamental representation $V$ emerged on the occasion of the proof of our earlier Lemma 7.10, which was not complete in version 7 of [11], and we worked on completion with the same methods. It was for us a sudden realization of an even greater recognition than we had realized up to that point of the contribution of the authors on extensions of Schur-Weyl duality which has historical interest on his own in representation theory, and more work by Wenzl and his collaborators, and several other authors in the classical and quantum cases for the various Lie types. We shall come back to this in more detail.
@@ 2. The main result
Let ${\mathfrak{g}}$ be a complex simple Lie algebra and let ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ be the module category of the corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ at level $k\in{\mathbb{N}}$, endowed with the ribbon braided tensor structure introduced by Huang and Lepowsky. Let $A(V_{{\mathfrak{g}}_{k}})$ be the Zhu algebra.
Let $V$ denote the fundamental representation of ${\mathfrak{g}}$ for each Lie type, defined in [91], (see also Sect. 11 for the classical types and $G_{2}$). We also denote by $V$ the corresponding quantized representation of $U_{q}({\mathfrak{g}})$, $q=e^{i\pi/\ell d}$, the associated object of the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$, with $\ell=k+\check{h}$, $\check{h}$ the dual Coxeter number of ${\mathfrak{g}}$. We use the same notation for the corresponding representation of the vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ and of the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$. Similarly for the irreducible objects $V_{\lambda}$ of the involved categories.
Let $A_{W}({\mathfrak{g}},q)$ denote the unitary compatible coboundary weak Hopf algebra associated to ${\mathcal{C}}({\mathfrak{g}},q)$ via the construction of a weak tensor structure on Wenzl functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$ following Theorems 30.17, in [11]. We shall give a brief exposition of $A_{W}({\mathfrak{g}},q)$ in Sect. 4.
@@
  • We have that:
@@
    The representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ admits a natural structure of ribbon braided tensor category obtained by transporting Huang-Lepowsky structure to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$. In this way, Zhu’s linear equivalence
@@
    Let $A(V_{{\mathfrak{g}}_{k}})$ be endowed with the twisted unitary compatible coboundary weak quasi-Hopf structure induced by the weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ via Wenzl path and the construction of the Drinfeld twist, following Theorem 31.1 in [11] (a), and let ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
    Then ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way with the structure derived from ${\mathcal{C}}({\mathfrak{g}},q)$ and the equivalence of Theorem 31.1 in [11],
    ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))\xrightarrow{}{\rm Rep}(A_{W}({\mathfrak{g}},q)).$
    Moreover the transport in (a) preserves the tensor products, the ribbon for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples $(V_{\lambda},V,V)$, $(V,V_{\lambda},V)$,$(V,V,V_{\lambda})$ with $V_{\lambda}$ an arbitrary irreducible of the fusion category.
@@
The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary compatible coboundary weak Hopf algebra, are given in Sects. 29, 30 in [11], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [91] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix. The former more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
@@
The construction of associativity morphisms satisfying the two hexagonal equations with respect to natural braiding morphisms associated to representations of the universal enveloping algebra $U({\mathfrak{g}})[[h]]$ of a simple Lie algebra (in a formal setting) with classical coproduct, were first solved by Drinfeld with the introduction and construction of quasi-Hopf algebras, via the Knizhnik-Zamolodchikov differential equation. This category is known as Drinfeld category, motivated by previous work by Kohno on monodromy of a connection associated to the vector representation for all ${\mathfrak{g}}$ of classical type [62]. Drinfeld developed the remarkable and beautiful Drinfeld-Kohno theorem connecting $U({\mathfrak{g}})[[h]]$ with its natural $R$-matrix and non-trivial co-associator with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ via a Drinfeld twist, describing a ribbon braided tensor equivalence. In this case, rigidity of Drinfeld category is derived from the equivalence constructed by Drinfeld in Drinfeld-Kohno theorem [23], [25], [26].
@@
In the same paper [62] Kohno had also considered monodromy of $n$-point functions of a conformal field theory on ${\mathbb{P}}^{1}$ associated to the affine Lie algebra $\hat{\mathfrak{s}l}_{2}$.
The mathematical contribution by Tsuchiya and Kanie [88], and the subsequent works by A. Wassermann [90] and Toledano-Laredo [87] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [91] that led to our approach. The reader may find reference to more of the several original papers in mathematics and physics in Sections 31–36 in [11].
@@
Similarities with the original Drinfeld-Kohno theorem also concern the identification method between the two associativity morphisms, in our case the twisted structure of that arising from quantum groups and that coming from Huang Lepowsky tensor structure both on the Zhu algebra. In Proposition 3.16 of [25], Drinfeld identifies the most difficult parts of the corresponding strictures for Drinfeld category using his uniqueness Theorem 3.15 based on cohomology of the same paper. In our case, we similarly develop a uniqueness theorem but for the associativity morphisms which we verify of our weak Hopf algebras $A_{W}({\mathfrak{g},q})$ and with very different methods, based on the generating representation $V$ for which we develop cohomological reduction Theorem 8.19 in [11] and extension to quantum case and different Lie types of Schur-Weyl duality. We discuss this more in the next two sections.
@@
The centralizer algebra of the fundamental representation for types $B$ and $D$ involve difficult work by Wenzl and his coauthors. Central theorems are Theorem 5.2 in [94] for the even case, and the previous Theorem 3.3 in [93] for the odd case. A special case has been studied by Rowell and Wenzl in [84].
@@ 13. Outlook
 [^1] Claims, arguments, and reasoning of the recent papers [6], [7], [8] are essentially dependent on the results and in-depth discussions between S.C. and C.P. that took place from 2018 to 2022 on the contents of an entire project including [11] and the topics of the above papers, during a period in which S.C. was a co-author, before breaking the collaboration in early 2023 by his decision, with a minor contribution in a draft of a single section. In addition, C.P. department digital devices were subject to a serious hacking attack for several months in 2021 during the previous collaboration, which was recognized by authorities, that severely damaged her intellectual property and health, and the hacked material has not been restituted.
Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of affine vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem. She is also grateful to Marco V. Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [11] and to Hans Wenzl for kind email correspondence on the centraliser algebras for the spinor representations, and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper, and to Makoto Yamashita for interesting discussions on [40] and [92] during his visit at the math department of Rome, Sapienza, and interesting comments on our work. Finally she thanks Pavel Etingof for informing her of [29].